# GST-UNet: Spatiotemporal Causal Inference with Time-Varying Confounders

### v1

@@ 3 Problem Formulation
Given $\mathbf{P}_{t}^{\tau}$, we show how to identify CAPOs from observational data. For $\tau\geq 2$, future covariates and outcomes (i.e. $\mathbf{X}_{t+1:t+\tau-1},\mathbf{Y}_{t+1:t+\tau-1}$) may influence subsequent treatments, creating time-varying confounding (Coston et al., 2020). These confounders, shaped by past treatments and outcomes, can alter future assignments and outcomes, forming feedback loops that simple “condition-on-history” adjustments miss. Hence, adjustment—e.g. via iterative G-computation—is needed to avoid bias. Figure 1 illustrates these complexities by comparing observational data (left) and hypothetical interventional data (right) for $\tau=2$. By contrast, when $\tau=1$, conditioning on $\mathbf{H}_{1:t}$ is sufficient under standard assumptions, as no future confounders lie between $\mathbf{A}_{t}$ and $\mathbf{Y}_{t+1}$. In other words, the following naive identification does not generally hold for $\tau>1$:
$\neq\mathbb{E}[\mathbf{Y}_{t+\tau}\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t},\mathbf{A}_{t:t+\tau-1}=\mathbf{a}_{t:t+\tau-1}]$ (3)
To address time-varying confounding, we adapt regression-based iterative G-computation (Bang and Robins, 2005; Robins and Hernan, 2008) to the spatiotemporal setting, providing a systematic way to adjust for evolving confounders and achieve unbiased CAPO estimates. We integrate our single time-series spatiotemporal setting with the G-computation framework in the following result:
Theorem 3.3 (Identification with G-Computation).
@@
Assume that 3.1 and 3.2 hold. Further, let $\mathbf{H}^{\mathbf{a}}_{1:t+k}:=(\mathbf{X}_{1:t+k},[\mathbf{A}_{1:t-1},\mathbf{a}_{t:t+k-1}],\mathbf{Y}_{1:t+k})$ denote the history where observed treatments from time $t$ onward are replaced by $\mathbf{a}_{t:t+k-1}$. Define recursively:
$Q_{\tau}(\mathbf{H}_{1:t+\tau-1},\mathbf{A}_{t+\tau-1})$
$\;\;=\mathbb{E}_{\mathbf{P}}[\mathbf{Y}_{t+\tau}\mid\phi(\mathbf{H}_{1:t+\tau-1},\mathbf{A}_{t+\tau-1})]$
$Q_{\tau-1}(\mathbf{H}_{1:t+\tau-2},\mathbf{A}_{t+\tau-2})$
$\;\;=\mathbb{E}_{\mathbf{P}}[Q_{\tau}(\mathbf{H}^{\mathbf{a}}_{1:t+\tau-1},\mathbf{a}_{t+\tau-1})\mid\phi(\mathbf{H}_{1:t+\tau-2},\mathbf{A}_{t+\tau-2})]$
@@
$Q_{1}(\mathbf{H}_{1:t},\mathbf{A}_{t})$
$\;\;=\mathbb{E}_{\mathbf{P}}[Q_{2}(\mathbf{H}^{\mathbf{a}}_{1:t+1},\mathbf{a}_{t+1})\mid\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})]$
Then:
$\mathbb{E}_{\mathbf{P}}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\phi(\mathbf{H}_{1:t},\mathbf{a}_{t})=\phi(\mathbf{h}_{1:t},\mathbf{a}_{t})]$
$=Q_{1}(\mathbf{h}_{1:t},\mathbf{a}_{t})$
We provide a proof of Theorem 3.3 in Appendix B. This result naturally motivates a recursive regression approach for spatiotemporal CAPO estimation, fitting each $Q_{k}(\cdot)$ in reverse order and substituting interventional treatments where required. In the next section, we detail how our end-to-end GST-UNet framework implements these ideas, providing a principled neural solution for time-varying confounding in spatiotemporal settings.
4 Methodology
Figure 2: Overview of the GST-UNet architecture. The embedding block (left) is a U-Net augmented with a ConvLSTM layer and attention gates. Its final feature map is passed to a set of G-heads (right), where each G-head $Q_{k}$ implements iterative G-computation (see Algorithm 1) to predict the potential outcomes or pseudo-outcomes at step $k$.
To address the challenges of time-varying confounders, spatial dependencies, and temporal carryover in observational data, we introduce GST-UNet, an end-to-end spatiotemporal neural model. Building on the identification guarantees in Theorem 3.3 and the single-series assumptions in 3.2, GST-UNet integrates iterative G-computation with a modular neural architecture to consistently estimate potential outcomes and treatment effects. We first convert the iterative G-computation result into a practical, recursive regression approach for CAPO estimation, before introducing the GST-UNet design in subsequent subsections.
4.1 Estimating CAPOs with Iterative G-Computation
While Theorem 3.3 motivates a recursive regression algorithm for each $Q_{k}$ ($k=1,\dots,\tau$), only $Q_{\tau}$ can be directly estimated from the prefix data. At the next step, $Q_{\tau-1}$ depends on $Q_{\tau}\bigl{(}\mathbf{H}^{\mathbf{a}}_{1:t+\tau-1},\mathbf{a}_{t+\tau-1}\bigr{)}$—where the observed treatments $\mathbf{A}_{t:t+\tau-1}$ are replaced by $\mathbf{a}_{t:t+\tau-1}$—but such substituted outcomes are not observed in the prefix data. Therefore, for $k<\tau$, we propose a procedure where we generate pseudo-outcomes by predicting with the previously learned $\widehat{Q}_{k+1}$. Going forward, we use $\widehat{F}$ to denote any quantity $F$ estimated from data. Formally, let $\phi\in\Phi$ be an embedding satisfying 3.2, and let $\mathcal{Q}$ be our function class for $Q_{k}$. We learn the sequence $\widehat{Q}_{\tau},\dots,\widehat{Q}_{1}$ from prefix data $\{\mathbf{P}_{t}^{\tau}:t=1,\dots,T-\tau\}$, via:
@@ 4.1 Estimating CAPOs with Iterative G-Computation
    Initialization. With the prefix data, fit $\widehat{Q}_{\tau}$ to predict $\mathbf{Y}_{t+\tau}$ from the embedding $\phi(\mathbf{H}_{1:t+\tau-1},\mathbf{A}_{t+\tau-1})$.
@@
      Generate pseudo-outcomes. Let $\widetilde{Y}_{t+k+1}\;=\;\widehat{Q}_{k+1}\bigl{(}\mathbf{H}^{\mathbf{a}}_{1:t+k},\mathbf{a}_{t+k}\bigr{)}$, where $\widehat{Q}_{k+1}$ is the previously learned function. These $\widetilde{Y}_{t+k+1}$ act as surrogates for $\mathbf{Y}_{t+k+1}$ in the prefix data.
@@
    $\widehat{\mathbb{E}}_{\mathbf{P}}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\phi(\mathbf{H}_{1:t},\mathbf{a}_{t})=\phi(\mathbf{h}_{1:t},\mathbf{a}_{t})]$
    $=\widehat{Q}_{1}\bigl{(}\mathbf{h}_{1:t},\,\mathbf{a}_{t}\bigr{)}.$
The iterative procedure yields a consistent estimator for the CAPOs if the $Q_{k}$’s are estimated consistently from data (Laan and Robins, 2003). In the following subsections, we instantiate this procedure in our GST-UNet architecture, illustrating how to incorporate spatial dependencies and interference into $\phi$ and each $Q_{k}$, and implement a streamlined, end-to-end training strategy that unifies history embeddings and outcome predictions.
@@ 4.2 Model Architecture
    Spatiotemporal Learning Module. A U-Net-based network augmented with ConvLSTM and attention gates for spatiotemporal processing.
@@
    Neural Causal Module. $\tau$ G-computation heads, each mapping the spatiotemporal features to the final outcome predictions in the iterative procedure.
@@
We illustrate the GST-UNet architecture in Figure 2 and describe its main components below.
@@
Spatiotemporal Learning Module. (1) Spatial Module. To efficiently process high-dimensional spatial data, we employ U-Net (Ronneberger et al., 2015), a fully convolutional architecture originally developed for biomedical image segmentation. It employs an encoder-decoder design with skip connections: the encoder progressively downsamples the spatial grid through convolution and pooling, while the decoder upsamples it back to the original resolution, merging encoder features at each scale. (2) Temporal Module. U-Net has limitations in capturing temporal information. To address this, we integrate a Convolutional Long Short-Term Memory (ConvLSTM) layer (Shi et al., 2015) to the U-Net encoder. This module captures temporal dependencies by maintaining a hidden state across time steps while aggregating spatial information through convolutions. After computing the final ConvLSTM state, we append static (time-invariant) covariates $\mathbf{V}$ as additional feature channels, ensuring the subsequent U-Net encoder-decoder has direct access to both temporal dynamics and static location-specific information. In the decoder, we incorporate attention gates (Oktay et al., 2018) to selectively highlight relevant spatial regions, refining skip connections and emphasizing critical global or local patterns. The embedding module ultimately produces a $d_{h}$-dimensional feature map of size $N_{X}\times N_{Y}$, capturing essential spatiotemporal context—including interference, spatial confounding, and static covariates—for downstream G-computation.
@@
Neural Causal Module. We attach $\tau$ G-computation heads to the U-Net’s final feature maps, corresponding to the $Q_{k}$ estimators in the iterative procedure (see Section 4.1). Each head can be a small convolutional module or a simple feed-forward network, depending on how much spatial structure remains to be captured. The information flow at the G-computation heads proceeds as follows: each head $Q_{k}$ ($k=1,\dots,\tau$) receives the $d_{h}\times N_{X}\times N_{Y}$ U-Net embedding $\widehat{\phi}\bigl{(}\mathbf{H}_{1:t+k-1},\mathbf{A}_{t+k-1}\bigr{)}$ (encompassing spatiotemporal and static context) and outputs an $N_{X}\times N_{Y}$ prediction for that time step. We refer to this as the supervision step, since $Q_{\tau}$ compares its predictions to the real observed outcomes $\mathbf{Y}_{t+\tau}$, anchoring the model in genuine data, while each $Q_{k<\tau}$ compares its predictions to pseudo-outcomes $\widetilde{\mathbf{Y}}_{t+k+1}$ provided by $\widehat{Q}_{k+1}$. These pseudo-outcomes arise in a subsequent generation step, wherein $Q_{k+1}$ processes the intervened history $\widehat{\phi}\bigl{(}\mathbf{H}^{\mathbf{a}}_{1:t+k},\mathbf{a}_{t+k}\bigr{)}$ in a detached forward pass (so $\widehat{Q}_{k+1}$ is not updated by $Q_{k}$’s loss), thereby creating surrogate targets for $Q_{k}$. This procedure realizes the iterative G-computation logic from Section 4.1, enabling GST-UNet to estimate future outcomes under various counterfactual treatments. By separating the spatiotemporal embedding from the G-heads, we maintain a common representation for all prefix data (see 3.2) and flexibly capture interference and spatial confounding. Each G-head enforces the proper temporal adjustments to yield bias-free counterfactual inference.
@@ 4.3 Training and Inference
1:  Input: Horizon $\tau$, prefix data $\{\mathbf{P}_{t}^{\tau}\}_{t=1}^{T-\tau}$, interventions $\mathbf{a}_{t:t+\tau-1}$, curriculum schedule $\alpha_{k}^{(e)}$, total epochs $E$.
@@
5:        (Supervision) For each prefix $i$, predict outcomes:
$\widehat{Y}_{t+k}^{(i)}=Q_{k}\bigl{(}\phi(\mathbf{H}_{1:t+k-1}^{(i)},\mathbf{A}_{t+k-1}^{(i)});\theta\bigr{)}.$
@@
$\widetilde{Y}_{t+k}^{(i)}=$
$\begin{cases}Q_{k}\Bigl{(}\phi\bigl{(}(\mathbf{H}^{\mathbf{a}}_{1:t+k-1})^{(i)},\,\mathbf{a}_{t+k-1}^{(i)}\bigr{)};\theta\Bigr{)},&k<\tau,\\ Y_{t+\tau}^{(i)},&k=\tau.\end{cases}$
@@
7:     end for
8:     (Loss aggregation) Compute the overall MSE loss
$\mathcal{L}(\theta;e)=\frac{1}{\tau}\sum_{k=1}^{\tau}\alpha_{k}^{(e)}\sum_{i}\bigl{(}\widehat{Y}_{t+k}^{(i)}-\widetilde{Y}_{t+k+1}^{(i)}\bigr{)}^{2}.$
9:     (Backward pass) Update $\theta$ by backpropagation.
10:  end for
11:  (Inference) Given a $\mathbf{h}_{1:t}$, return $Q_{1}(\phi(\mathbf{h}_{1:t},\mathbf{a}_{t});\widehat{\theta})$.
A key obstacle in learning from a single spatiotemporal series is that we must splice the data into many prefixes (3.2), then estimate all $Q_{k}$ in an iterative G-computation procedure (Section 4.1). In principle, we could train each G-head $Q_{k}$ sequentially (from $\tau$ down to $1$) by feeding its pseudo-outcomes to the next head. However, this creates complications when sharing the U-Net embedding $\phi$ across all heads: each $Q_{k}$ might attempt to tailor $\phi$ to its own objective, leading to misaligned training signals.
Joint Loss and Multi-Task Training. To address this issue, we employ a joint (or multi-task) training approach (Caruana, 1997; Evgeniou and Pontil, 2004) by aggregating the loss terms from all G-heads into a single objective, then backpropagating once per batch. Concretely, for each head $Q_{k}$, let $\widetilde{Y}_{t+k+1}$ be the real outcomes if $k=\tau$ or pseudo-outcomes (generated by $\widehat{Q}_{k+1}$) if $k<\tau$. Our head-specific loss is a mean squared error (MSE) over all prefix samples:
@@
where $\theta$ encompasses all model parameters (the shared U-Net embedding $\phi$ and the G-heads $Q_{k}$). Let $\alpha_{k}^{(e)}$ denote a head-weight for epoch $e$. We then form the overall training objective at epoch $e$ by
@@
Curriculum Training. A naive implementation of (4) – where we give each G-head an equal weights – can be suboptimal: early in training, $Q_{\tau}$ (which sees real data) is inaccurate, and thus the pseudo-outcomes generated for $Q_{k<\tau}$ are effectively noise. Consequently, heads $Q_{1},\dots,Q_{\tau-1}$ may overfit to poor targets before $Q_{\tau}$ has converged, leading to suboptimal solutions. To mitigate this, we employ a curriculum training approach (Bengio et al., 2009), gradually increasing the loss weight of earlier heads as $Q_{\tau}$ improves.
@@
Hence, during epochs $1\leq e\leq e_{c}$ (phase $p(e)=1$), only $Q_{\tau}$ is active with $\alpha_{\tau}^{(e)}=1$; in the next interval $e_{c}<e\leq 2e_{c}$ (phase $p(e)=2$), $Q_{\tau}$ and $Q_{\tau-1}$ each have weight $1/2$, and so on, until phase $\tau$ sets all heads to uniform weight $1/\tau$. For $e>\tau\,e_{c}$, we continue training with $\alpha_{k}^{(e)}=1/\tau$ across all G-heads for additional joint refinement. This progressive schedule ensures $Q_{\tau}$ becomes reasonably accurate on the observed data before earlier heads rely on its pseudo-outcomes. The single hyperparameter $e_{c}$ succinctly controls how quickly we introduce each head, preventing excessive noise in early training.
We also adopt standard neural network practices, including mini-batch optimization and early stopping, to stabilize training and mitigate overfitting. At inference time, given a new history $\mathbf{h}_{1:t}$ and an interventional sequence $\mathbf{a}_{t:t+\tau-1}$, we compute $\widehat{Q}_{1}(\phi(\mathbf{h}_{1:t},\mathbf{a}_{t});\theta)$ as our target CAPO estimate. We sketch the overall training and inference procedure in Algorithm 1.

### v2

@@ 4.1 Identification via Representation-Based G-Computation
Given $\mathbf{P}_{t}^{\tau}$, we next show how to identify CAPOs from observational data. For horizons $\tau\geq 2$, future covariates and outcomes (i.e. $\mathbf{X}_{t+1:t+\tau-1},\mathbf{Y}_{t+1:t+\tau-1}$) can influence subsequent treatments, inducing time-varying confounding [13]. Such feedback violates standard "condition-on-history" adjustments and leads to biased estimates. Figure˜1 illustrates these dependencies by contrasting observational data (left) and hypothetical interventions (right) for $\tau=2$. By contrast, when $\tau=1$, conditioning on $\mathbf{H}_{1:t}$ is sufficient under standard assumptions, as no future confounders intervene between $\mathbf{A}_{t}$ and $\mathbf{Y}_{t+1}$. Formally, the following naive identification fails to hold for $\tau>1$:
$\mathbb{E}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t}]\neq\mathbb{E}[\mathbf{Y}_{t+\tau}\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t},\mathbf{A}_{t:t+\tau-1}=\mathbf{a}_{t:t+\tau-1}]$ (2)
To correct this bias, we adapt regression-based iterative G-computation [3, 35] to the spatiotemporal setting, yielding a principled adjustment procedure for evolving confounders and valid CAPO estimation. We formalize this connection in the following result:
Theorem 1 (Identification with G-Computation).
@@
Assume that Assumption˜1 and Assumption˜2 hold. Further, let $\mathbf{H}^{\mathbf{a}}_{1:t+k}:=(\mathbf{X}_{1:t+k},[\mathbf{A}_{1:t-1},\mathbf{a}_{t:t+k-1}],\mathbf{Y}_{1:t+k})$ denote the history where observed treatments from time $t$ onward are replaced by $\mathbf{a}_{t:t+k-1}$. Define recursively:
$Q_{\tau}(\mathbf{H}_{1:t+\tau-1},\mathbf{A}_{t+\tau-1})=\mathbb{E}_{\mathbf{P}}[\mathbf{Y}_{t+\tau}\mid\phi(\mathbf{H}_{1:t+\tau-1},\mathbf{A}_{t+\tau-1})]$
$Q_{\tau-1}(\mathbf{H}_{1:t+\tau-2},\mathbf{A}_{t+\tau-2})=\mathbb{E}_{\mathbf{P}}[Q_{\tau}(\mathbf{H}^{\mathbf{a}}_{1:t+\tau-1},\mathbf{a}_{t+\tau-1})\mid\phi(\mathbf{H}_{1:t+\tau-2},\mathbf{A}_{t+\tau-2})]$
@@
$Q_{1}(\mathbf{H}_{1:t},\mathbf{A}_{t})=\mathbb{E}_{\mathbf{P}}[Q_{2}(\mathbf{H}^{\mathbf{a}}_{1:t+1},\mathbf{a}_{t+1})\mid\phi(\mathbf{H}_{1:t},\mathbf{A}_{t})]$
Then $\mathbb{E}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\mathbf{H}_{1:t}=\mathbf{h}_{1:t}]=Q_{1}(\mathbf{h}_{1:t},\mathbf{a}_{t})$.
We provide a proof of Theorem˜1 in Appendix˜B. This result naturally motivates a recursive regression approach for spatiotemporal CAPO estimation, fitting each $Q_{k}(\cdot)$ in reverse order and substituting interventional treatments where required.
4.2 Estimation via Iterative G-Computation
While Theorem˜1 motivates a recursive regression algorithm for each $Q_{k}$ ($k=1,\dots,\tau$), only $Q_{\tau}$ can be directly estimated from the prefix data. At the next step, $Q_{\tau-1}$ depends on $Q_{\tau}\bigl(\mathbf{H}^{\mathbf{a}}_{1:t+\tau-1},\mathbf{a}_{t+\tau-1}\bigr)$—where the observed treatments $\mathbf{A}_{t:t+\tau-1}$ are replaced by $\mathbf{a}_{t:t+\tau-1}$—but such substituted outcomes are not observed in the prefix data. Therefore, for $k<\tau$, we propose a procedure where we generate pseudo-outcomes by predicting with the previously learned $\widehat{Q}_{k+1}$. Going forward, we use $\widehat{F}$ to denote any quantity $F$ estimated from data. Formally, let $\phi\in\Phi$ be an embedding satisfying Assumption˜2, and let $\mathcal{Q}$ be our function class for $Q_{k}$. We learn the sequence $\widehat{Q}_{\tau},\dots,\widehat{Q}_{1}$ from prefix data $\{\mathbf{P}_{t}^{\tau}:t=1,\dots,T-\tau\}$, via:
@@ 4.2 Estimation via Iterative G-Computation
    Initialization. Fit $\widehat{Q}_{\tau}$ to predict $\mathbf{Y}_{t+\tau}$ from the prefix embedding $\phi(\mathbf{H}_{1:t+\tau-1},\mathbf{A}_{t+\tau-1})$.
@@
      Generate pseudo-outcomes. Let $\widetilde{Y}_{t+k+1}=\widehat{Q}_{k+1}\bigl(\mathbf{H}^{\mathbf{a}}_{1:t+k},\mathbf{a}_{t+k}\bigr)$, where $\widehat{Q}_{k+1}$ was learned in the previous step. These $\widetilde{Y}_{t+k+1}$ act as surrogates for $\mathbf{Y}_{t+k+1}$ in the prefix data.
@@
    $\widehat{\mathbb{E}}_{\mathbf{P}}[\mathbf{Y}_{t+\tau}[\mathbf{a}_{t:t+\tau-1}]\mid\phi(\mathbf{H}_{1:t},\mathbf{a}_{t})=\phi(\mathbf{h}_{1:t},\mathbf{a}_{t})]=\widehat{Q}_{1}\bigl(\mathbf{h}_{1:t},\,\mathbf{a}_{t}\bigr).$
The iterative regression procedure yields consistent CAPO estimates provided each stage $Q_{k}$ is estimated consistently from data [22]. Informally, if the learned embedding $\widehat{\phi}$ converges to the true time-invariant representation $\phi$, and small perturbations in $\phi$ or $\widehat{Q}_{k}$ lead to proportionally small changes in predictions, then the overall recursive estimator remains consistent. These regularity conditions—formalized through uniform stochastic equicontinuity—are detailed in Appendix˜C. Formally, we state the following theorem:
Theorem 2 (Consistency of Iterative G-Computation in Spatiotemporal Settings).
Assume Assumptions˜1 and 2 and that (a) the learned embedding $\widehat{\phi}$ is $L_{2}$-consistent for $\phi$, and (b) each regression head $\widehat{Q}_{k}$ consistently estimates $Q_{k}$ and is uniformly well-behaved [^1] We formalize ”well-behaved” via uniform stochastic equicontinuity and continuity in Appendix C. on $\mathrm{Im}\,\phi$ (intuitively, small input perturbations induce small output changes). Let $\mathbf{Z}_{k}:=(\mathbf{H}_{1:t+k},\mathbf{A}_{t+k})$ denote the history–action pair at step $k$. Then
$\bigl\|\widehat{Q}_{1}(\mathbf{Z}_{0};\widehat{\phi})-Q_{1}(\mathbf{Z}_{0};\phi)\bigr\|_{2}=o_{p}(1),$
so the recursive estimator $\widehat{Q}_{1}$ of the CAPO is probabilistically consistent.
We provide a proof of Theorem˜2 in Appendix˜C. In the following section, we instantiate this procedure in our GST-UNet architecture, illustrating how to incorporate spatial dependencies and interference into $\phi$ and each $Q_{k}$, and implement a streamlined, end-to-end training strategy that unifies history embeddings and outcome predictions.
5 GST-UNet Implementation
Figure 2: Overview of the GST-UNet architecture. The spatiotemporal learning module (left) is a U-Net augmented with a ConvLSTM layer and attention gates. Its final feature map is passed to a set of G-heads (right), where each G-head $Q_{k}$ implements iterative G-computation (see Algorithm˜1).
The theoretical results above establish how CAPOs can be identified and consistently estimated from a single spatiotemporal trajectory. We now provide a concrete neural implementation of this procedure. GST-UNet instantiates the iterative G-computation framework with a spatiotemporal deep architecture that embeds strong inductive biases—locality, translation invariance, and temporal smoothness—well suited to data-scarce settings. While alternative backbones could be employed, our U-Net with ConvLSTM and attention offers a natural choice for learning stable, history-invariant representations that satisfy Assumption˜2. We now describe the architecture and the training procedure that realizes the GST-UNet (Algorithm˜1).
@@ 5.1 Model Architecture
    Spatiotemporal Learning Module: a U-Net-based network augmented with ConvLSTM and attention gates for spatiotemporal processing.
@@
    Neural Causal Module: $\tau$ G-computation heads, each mapping the spatiotemporal features to the final outcome predictions in the iterative procedure.
@@
We illustrate the GST-UNet architecture in Figure˜2 and describe its main components below.
@@
Spatiotemporal Learning Module. (1) Spatial Module. While our framework is agnostic to the choice of spatiotemporal learning module, we adopt a U-Net with ConvLSTM and attention due to its strong performance in data-scarce regimes. To efficiently process high-dimensional spatial data, we employ U-Net [37], a fully convolutional architecture originally developed for biomedical image segmentation. It employs an encoder-decoder design with skip connections: the encoder progressively downsamples the spatial grid through convolution and pooling, while the decoder upsamples it back to the original resolution, merging encoder features at each scale. (2) Temporal Module. U-Net has limitations in capturing temporal information. To address this, we integrate a Convolutional Long Short-Term Memory (ConvLSTM) layer [40] to the U-Net encoder. This module captures temporal dependencies by maintaining a hidden state across time steps while aggregating spatial information through convolutions. After computing the final ConvLSTM state, we append static (time-invariant) covariates $\mathbf{V}$ as additional feature channels, ensuring the subsequent U-Net encoder-decoder has direct access to both temporal dynamics and static location-specific information. In the decoder, we incorporate attention gates [29] to selectively highlight relevant spatial regions, refining skip connections and emphasizing critical global or local patterns. The embedding module ultimately produces a $d_{h}$-dimensional feature map of size $N_{X}\times N_{Y}$, capturing essential spatiotemporal context—including interference, spatial confounding, and static covariates—for downstream G-computation.
@@
Neural Causal Module. We attach $\tau$ G-computation heads to the U-Net’s final feature maps, corresponding to the $Q_{k}$ estimators in the iterative procedure (see Section˜4.2). Each head can be a small convolutional module or a simple feed-forward network, depending on how much spatial structure remains to be captured. The information flow at the G-computation heads proceeds as follows: each head $Q_{k}$ ($k=1,\dots,\tau$) receives the $d_{h}\times N_{X}\times N_{Y}$ U-Net embedding $\widehat{\phi}\bigl(\mathbf{H}_{1:t+k-1},\mathbf{A}_{t+k-1}\bigr)$ (encompassing spatiotemporal and static context) and outputs an $N_{X}\times N_{Y}$ prediction for that time step. We refer to this as the supervision step, since $Q_{\tau}$ compares its predictions to the real observed outcomes $\mathbf{Y}_{t+\tau}$, anchoring the model in genuine data, while each $Q_{k<\tau}$ compares its predictions to pseudo-outcomes $\widetilde{\mathbf{Y}}_{t+k+1}$ provided by $\widehat{Q}_{k+1}$. These pseudo-outcomes arise in a subsequent generation step, wherein $Q_{k+1}$ processes the intervened history $\widehat{\phi}\bigl(\mathbf{H}^{\mathbf{a}}_{1:t+k},\mathbf{a}_{t+k}\bigr)$ in a detached forward pass (so $\widehat{Q}_{k+1}$ is not updated by $Q_{k}$’s loss), thereby creating surrogate targets for $Q_{k}$. This procedure realizes the iterative G-computation logic from Section˜4.2, enabling GST-UNet to estimate future outcomes under various counterfactual treatments. By separating the spatiotemporal embedding from the G-heads, we maintain a common representation for all prefix data (see Assumption˜2) and flexibly capture interference and spatial confounding. Each G-head enforces the proper temporal adjustments to yield bias-free counterfactual inference.
@@ 5.2 Training and Inference
1:Input: Horizon $\tau$, prefixes $\{\mathbf{P}_{t}^{\tau}\}_{t=1}^{T-\tau}$, interventions $\mathbf{a}_{t:t+\tau-1}$, curriculum $\alpha_{k}^{(e)}$, total epochs $E$.
@@
5:   (Supervision) For each prefix $i$, predict outcomes $\widehat{Y}_{t+k}^{(i)}=Q_{k}\bigl(\phi(\mathbf{H}_{1:t+k-1}^{(i)},\mathbf{A}_{t+k-1}^{(i)});\theta\bigr)$.
@@
$\widetilde{Y}_{t+k}^{(i)}=\begin{cases}Q_{k}\Bigl(\phi\bigl((\mathbf{H}^{\mathbf{a}}_{1:t+k-1})^{(i)},\,\mathbf{a}_{t+k-1}^{(i)}\bigr);\theta\Bigr),&k<\tau,\\ Y_{t+\tau}^{(i)},&k=\tau.\end{cases}$
@@
7:  (Loss aggregation) Compute the MSE loss $\mathcal{L}(\theta;e)=\frac{1}{\tau}\sum_{k=1}^{\tau}\alpha_{k}^{(e)}\sum_{i}\bigl(\widehat{Y}_{t+k}^{(i)}-\widetilde{Y}_{t+k+1}^{(i)}\bigr)^{2}$.
8:  (Backward pass) Update $\theta$ by backpropagation.
[…]
While each G-head $Q_{k}$ could be trained sequentially–from $Q_{\tau}$ down to $Q_{1}$–by passing pseudo-outcomes backward through time, this creates a conflict when all heads share the same U-Net embedding $\phi$. Specifically, each $Q_{k}$ may push $\phi$ toward optimizing its own objective, resulting in misaligned training signals and unstable learning.
Joint Loss and Multi-Task Training. To address this issue, we employ a joint (or multi-task) training approach [9, 15] by aggregating the loss terms from all G-heads into a single objective, then backpropagating once per batch. Concretely, for each head $Q_{k}$, let $\widetilde{Y}_{t+k+1}$ be the real outcomes if $k=\tau$ or pseudo-outcomes (generated by $\widehat{Q}_{k+1}$) if $k<\tau$. Our head-specific loss is a mean squared error (MSE) over all prefix samples:
@@
where $\theta$ encompasses all model parameters (the shared U-Net embedding $\phi$ and the G-heads $Q_{k}$).
Let $\alpha_{k}^{(e)}$ denote a head-weight for epoch $e$. We then form the overall training objective at epoch $e$ by
@@
Curriculum Training. A naive implementation of Eq. (3)–where each G-head is given equal weight–can be suboptimal: early in training, $Q_{\tau}$ (which sees real data) is inaccurate, and the pseudo-outcomes generated for $Q_{k<\tau}$ are effectively noise. Consequently, $Q_{1},\dots,Q_{\tau-1}$ may overfit to poor targets before $Q_{\tau}$ has converged, leading to suboptimal solutions. To mitigate this, we employ a curriculum training approach [5], gradually increasing the loss weight of earlier heads as $Q_{\tau}$ improves.
@@
Hence, during epochs $1\leq e\leq e_{c}$ (phase $p(e)=1$), only $Q_{\tau}$ is active with $\alpha_{\tau}^{(e)}=1$; in the next interval $e_{c}<e\leq 2e_{c}$ (phase $p(e)=2$), $Q_{\tau}$ and $Q_{\tau-1}$ each have weight $1/2$, and so on until all heads are active with uniform weight $1/\tau$. For $e>\tau e_{c}$, training continues with $\alpha_{k}^{(e)}=1/\tau$ for all heads. This schedule ensures $Q_{\tau}$ becomes reasonably accurate before earlier heads rely on its pseudo-outcomes. The hyperparameter $e_{c}$ controls the pacing, helping prevent early training noise.
We also adopt standard neural network practices, including mini-batch optimization and early stopping, to stabilize training and mitigate overfitting. At inference time, given a new history $\mathbf{h}_{1:t}$ and an interventional sequence $\mathbf{a}_{t:t+\tau-1}$, we compute $\widehat{Q}_{1}(\phi(\mathbf{h}_{1:t},\mathbf{a}_{t});\theta)$ as our target CAPO estimate. We sketch the overall training and inference procedure in Algorithm˜1.